# Decentralized Cooperation in Heterogeneous Multi-Agent Reinforcement Learning via Graph Neural Network-Based Intrinsic Motivation

### v1

@@ Abstract
Introduction
@@
The paradigm of Multi-agent Reinforcement Learning (MARL) is rapidly emerging to be pivotal in a broad spectrum of practical applications such as resource management (Ying and Dayong 2005), autonomous vehicles (Cao et al. 2013), traffic signal control (Calvo and Dusparic 2018a), supply chain management (Fuji et al. 2018), robotics (Lillicrap et al. 2019), robot swarms (Hüttenrauch, Šošić, and Neumann 2017), etc. These applications generally benefit from the efficient use of the diverse capabilities of heterogeneous agents. Moreover, the successful execution of tasks in these multi-agent systems requires the agents to adapt their behaviors to other agents for effective coordination rather than operating independently. The real-world deployment of these MARL systems typically involves the agents relying solely on the local environmental information and learning policies with infrequent environmental rewards (Wiewiora 2010; Wakilpoor et al. 2020).
@@
Applications such as package transport (Gerkey and Mataric 2002), traffic lights control (Calvo and Dusparic 2018b), disaster response (Michael et al. 2012), agriculture (Ju and Son 2019), etc. utilize agent heterogeneity such as distinct physical and behavioral traits of agents. Heterogeneity is also vital in multi-robot tasks as it enables efficient characterization and discovery of diverse behaviors, improving learning performance (Liu et al. 2022). On the other hand, the dependency on reward signals for the agent’s learning process introduces the issue of reward sparsity (Hare 2019). Due to the lack of frequent feedback signals from the environment and the non-trivial nature of manually designing reward functions, MARL systems need to be robust enough to deal with infrequent environmental rewards.
@@
Furthermore, in addition to these challenges, the majority of real-world applications constrain the agents to act in a decentralized manner, and under partial observability, where each agent has a partial view of the shared environment. As a result, it is impractical for them to learn cooperative behaviors by utilizing a centralized algorithm that possesses global knowledge of all the agents and the state space (Iqbal and Sha 2019; Liu, Yeh, and Schwing 2019). In comparison to centralized training and full observability, the challenges of agent heterogeneity and reward sparsity are more severe in decentralized training under partial observability (Ma et al. 2022). Despite the aforementioned requirements for real-world deployments, most of the existing solutions to date either utilize global parameter sharing among all the agents or employ a centralized critic with global knowledge.
@@
To the best of our knowledge, no prior research has addressed the issue of cooperative heterogeneous MARL in decentralized training settings under the practical constraints of real-world applications, such as partial observability and reward sparsity (see Section Related Works for more details). We propose CoHet, an algorithm that facilitates heterogeneous agent cooperation addressing the constraints required for real-world deployments. CoHet does not require any prior knowledge of agent heterogeneity (e.g speed, size, type, agent index). It employs an architecture for learning heterogeneous MARL policies by utilizing a novel Graph Neural Network-based intrinsic motivation/reward calculation mechanism. In summary, our specific contributions are as follows:
@@
A Novel Intrinsic Reward Mechanism: We introduce a novel self-supervised intrinsic reward calculation algorithm - CoHet - utilizing the underlying communication graph of a Graph Neural Network (GNN). In comparison to previous methods, CoHet accurately estimates the intrinsic rewards in the presence of agent heterogeneity (e.g. physical attributes/composition, behavioral differences) by using only the agents’ local neighborhood information and their dynamics model predictions instead of relying on a centralized critic or external controller. We present two formulations of the CoHet algorithm, one that utilizes the neighborhood predictions (CoHetteam), and another that uses the agent’s own predictions (CoHetself) for intrinsic reward calculation.
@@
Integration with Established Algorithms: Our standalone intrinsic motivation architecture can be integrated with existing decentralized heterogeneous policy learning algorithms, thus enhancing performance in cooperative MARL benchmarks. We demonstrate thisCodebase - github.com/jahirsadik/CoHet-Implementation by incorporating the state-of-the-art HetGPPO algorithm (Bettini, Shankar, and Prorok 2023a), leveraging its underlying GNN communication graph for intrinsic reward calculation. In contrast to previous heterogeneous policy learning techniques, this formulation requires no prior knowledge of the types of agent heterogeneity, indexing, etc.
@@
Extensive Validation and Scalability: We validate CoHet in the presence of heterogeneous agents in six different scenarios in the Multi-agent Particle Environment (MPE) and Vectorized Multi-Agent Simulator (VMAS) benchmarks, showing superior performance to the baseline HetGPPO across all scenarios. CoHet also outperforms the state-of-the-art MARL policy learning algorithm, Independent Proximal Policy Optimization (IPPO), in four of the six scenarios. We present findings on the impact of agent dynamics models on the intrinsic reward calculation, compare the two variants of the algorithm, and demonstrate its robustness to an increasing number of heterogeneous agents in a shared environment.
The subsequent sections will establish the necessary foundation of our work and explore our contributions in detail.
Related Works
@@
In response to the challenges posed by agent heterogeneity and the sparsity of rewards, most existing literature addresses one or the other. A prevalent method for emulating heterogeneous behavior involves augmenting each agent’s observation space with a unique index that represents the agent’s type (Foerster et al. 2016; Gupta, Egorov, and Kochenderfer 2017). These methods necessitate prior knowledge of the types of agent heterogeneity. Most of the literature on heterogeneous MARL focuses on tackling the problem of behavioral heterogeneity for physically identical agents (Wang et al. 2020a, b; Li et al. 2021; Lowe et al. 2020). They cannot be applied to all types of heterogeneity as they suggest solutions that are specific to problems within specific sub-classes of heterogeneity. Moreover, their use in partially observable systems is limited due to the lack of inter-agent communication. A notable study on heterogeneous MARL (Bettini, Shankar, and Prorok 2023a), proposes an algorithm termed HetGPPO (Heterogeneous GNN-based Proximal Policy Optimization), that is capable of learning heterogeneous decentralized policies in partially observable scenarios. It uses a GNN-based communication layer for sharing information among agents within local neighborhoods, thereby mitigating the effects of partial observability. Unlike previous methods, their reliance on only the local information enables Decentralized Training with Decentralized Execution (DTDE). Similar to their work, CoHet utilizes only the local neighborhood information under the constraints of DTDE.
@@
Although the aforementioned studies address the problem of agent heterogeneity, typically involving centralized training settings, parameter sharing or prior knowledge of agent heterogeneity, they do not consider the issues with reward sparsity. For tackling the sparse-reward problem in MARL, learning with the efficient exploration of the state space is a widely adopted approach (Liu et al. 2021; Mahajan et al. 2020; Christianos, Schäfer, and Albrecht 2021). Dense scalar reward signals termed “Intrinsic Motivation”, are often used to encourage exploration or coordination among agents (Du et al. 2019; Wu et al. 2021). Studies with a primary focus on cooperative tasks rely on the ability of agents to represent the dynamics of other agents in a shared environment to calculate these intrinsic rewards (Jaques et al. 2019; Stadie, Levine, and Abbeel 2015; Jeon et al. 2022). In case of the algorithm proposed by Ma et al. (2022), termed ELIGN (Expectation Alignment as a Multi-Agent Intrinsic Reward), the intrinsic rewards foster inter-agent coordination in a decentralized manner. However, a significant drawback of using an agent’s own dynamics model as a proxy to calculate neighborhood predictions in ELIGN is that it becomes more challenging for agents to accurately model the dynamics of other agents in the presence of agent heterogeneity. Inaccuracies in the dynamics model can result in misleading alignment signals, and as a result, ELIGN scales poorly in the presence of heterogeneous agents. In contrast, CoHet utilizes the local neighborhood information passed via the underlying GNN-communication graph, to more accurately model the heterogeneity among agents.
@@
There are only a couple of existing research works that simultaneously tackle agent heterogeneity under constraints of partial observability and reward sparsity (Zheng et al. 2020; Andres, Villar-Rodriguez, and Ser 2022). In the former work, heterogeneity is defined differently, referring to a mixture of on-policy, off-policy, and Evolutionary Algorithm (EA) agents and not the diverse physical and behavioral traits of agents. Moreover, their adoption of a local-global memory replay prevents them from undergoing training in a fully decentralized manner. The latter leverages intrinsic motivation in order to tackle reward sparsity similar to our work. They also address collaboration among heterogeneous agents. However, their utilization of a centralized critic that merges all agent parameters into a single network constrains its applicability in the DTDE paradigm. Despite its significance, the lack of solutions in decentralized training settings limits the deployment of MARL agents in practical applications. CoHet addresses this notable research gap in the area of cooperative heterogeneous MARL, by fostering cooperation in a decentralized manner, under the real-world challenges of partial observability and reward sparsity.
Background
@@
In this section, we discuss the Markov games framework, which we use to formulate our problem, and the message-passing Graph Neural Network, which is used for differentiable inter-agent communication of local observations and predictions among agents.
Markov Games
@@
The Markov games framework (Littman 1994) is a generalization of a Markov decision process to the case of multiple agents with cooperating or competing goals. The Partially Observable Markov Games (POMG) framework is used under conditions of partial observability. It is defined using the tuple —
@@
where,
  • $\mathcal{V}=\{1,2,\ldots,N\}$ is the set of all $N$ agents
  • $\mathcal{S}$ is the entire state space shared among all the agents
  • $\mathcal{O}\equiv\mathcal{O}_{1}\times\mathcal{O}_{2}\times\ldots\times\mathcal{O}_{n}$ is the observation space of all the agents, where $\mathcal{O}_{i}\subseteq\mathcal{S}$ and $\forall i\in\mathcal{V}$
  • $\mathcal{A}\equiv\mathcal{A}_{1}\times\mathcal{A}_{2}\times\ldots\times\mathcal{A}_{n}$ is the action space of all the agents
  • $o_{i}\in\mathcal{O}_{i}$ is the observation instance of agent $i\in\mathcal{V}$, which is a partial view of the state space $\mathcal{S}$
  • $\mathcal{R}_{i}\colon\mathcal{S}\times\mathcal{A}\times\mathcal{S}\mapsto\mathbb{R}$ is the reward function for agent $i$, where $i\in\mathcal{V}$
  • $\mathcal{T}\colon\mathcal{S}\times\mathcal{A}\times\mathcal{S}\mapsto[0,1]$ is the stochastic state transition model, where each agent transitions to the next state, given the current state and the action taken
  • $\gamma$ is the discount factor such that $0\leq\gamma\leq 1$
Additionally, the reward that agent $i$ receives at time step $t$ can be denoted using $r_{i}^{t}=\mathcal{R}_{i}(s^{t},a^{t},s^{t+1})$, where $a^{t}=(a_{1}^{t},a_{2}^{t},\ldots,a_{n}^{t})$ is the joint action taken by all the agents at time step $t$. The goal of each agent is to maximize the total expected discounted return $R_{i}=\sum^{T}_{t=0}\gamma^{t}r_{i}^{t}$ over the course of an episode with horizon $T$.
Message-passing Graph Neural Network
The message-passing technique (Kipf and Welling 2017) used by GNNs to transfer information from one node to another has proven to be an effective learning framework for understanding the patterns in a graph, the neighborhood of nodes, and the sub-graphs in large graphs (Zhou et al. 2021). In our problem formulation, given a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$,
@@
The use of only the non-absolute observation features as agent embeddings allows the outputs of the message-passing GNN kernel to be invariant to geometric translations, thereby enhancing generalization. For the message-passing, at each time step, the agent embeddings and edge features are first computed and then utilized in the message-passing GNN kernel to learn the local sub-graph. Information is iteratively passed between adjacent nodes along the edges of the graph structure. In order to incorporate the CoHet architecture on top of the underlying GNN formulation of HetGPPO, the inputs to our message-passing GNN kernel consist of the agent embedding $z_{i}=\omega_{\theta_{i}}(x_{i})$ and edge attributes $e_{ij}=p_{ij}\|v_{ij}$. Here, $\omega_{\theta_{i}}$ represents a Multi-Layer Perceptron (MLP) encoder with parameters $\theta_{i}$. Using local information from all neighbors $j\in\mathcal{N}_{i}$, the GNN model output $h_{i}$ for agent $i$ is calculated in Equation 1.
@@
In Equation 1, $\psi_{\theta_{i}}$ and $\phi_{\theta_{i}}$ are two MLPs parameterized by $\theta_{i}$ and the aggregation operator $\bigoplus$ sums the $\phi_{\theta_{i}}$ outputs for all the neighbors of agent $i$. Finally, two distinct MLP decoders take the GNN output $h_{i}$ and produce the value $V_{i}(o_{\mathcal{N}_{i}})$ and the action $a_{i}$, distributed according to $a_{i}\sim\pi_{i}(\cdot\mid o_{\mathcal{N}_{i}})$ This formulation of GNN, allows us to utilize it for both intrinsic reward calculation and heterogeneous policy learning of HetGPPO, based on local neighborhood information.
The CoHet Algorithm
In this section, we introduce CoHet, a decentralized algorithm designed to enhance cooperation among heterogeneous agents in partially observable environments with sparse rewards. It provides a standalone self-supervised intrinsic reward architecture that can be incorporated with existing decentralized policy optimization algorithms. It fosters the learning of collaborative behaviors by reducing future uncertainty within each agent’s neighborhood. CoHet encourages the agents to align their actions with their neighbors’ predictions by imposing intrinsic reward penalties that deter deviations from such alignment. Furthermore, these calculated rewards serve as a source of dense reward signals that facilitate policy learning in numerous real-world tasks where manually designing reward functions is infeasible.
Algorithm Description
@@
29:         $r^{t}_{total}\leftarrow r^{t}_{total_{1}}\|r^{t}_{total_{2}}\|\ldots\|r^{t}_{total_{N}}$
30:         $\mathcal{D}_{k}\leftarrow\mathcal{D}_{k}\cup(o^{t},a^{t},r_{total}^{t},o^{t+1})$
@@
In Algorithm 1, we start by initializing the model for each agent which includes the encoder $\omega_{i}$, two multi-layer perceptrons (MLPs) $\psi_{i}$ and $\phi_{i}$, $\pi$-decoder $\Omega_{i}$, value decoder $\Gamma_{i}$, and dynamics model $f_{i}$. At each training iteration, we begin by collecting the observations $o_{i}^{t}$. We trim the position $p_{i}^{t}$ and velocity $v_{i}^{t}$ from the observation to obtain the non-absolute features $x_{i}^{t}$ of each agent observation in Line 7, and use that to calculate the node embedding $z_{i}^{t}$ in Line 8. The position $p_{i}^{t}$ and velocity $v_{i}^{t}$ from the observations are used as edge embeddings in Line 11, this allows for GNN outputs to be invariant to geometric translations. Neighboring agent embeddings $\{z_{j}^{t}\mid j\in\mathcal{N}_{i}^{t}\}$ are obtained through the GNNs underlying differentiable communication channel. These neighborhood embeddings, along with edge features are aggregated to calculate the GNN model output in Line 12, which is then decoded to produce action and value outputs in Line 14 and Line 15. The joint action $a^{t}$, of all the agents, is formed by concatenating individual agent actions at time $t$, in Line 17. Individual agent actions $a^{t}_{i}$, are later used for intrinsic reward calculation, and the concatenated actions of all agents $a^{t}$ are used for policy optimization.
The procedure for policy optimization as discussed above is visualized in Figure 1(b), where the reward assigned to the agents at each time step is augmented with the dense intrinsic reward calculated via the process outlined in Figure 1(a). Prior intrinsic motivation-based solutions have been shown to alleviate the effects of reward sparsity (as discussed in the ‘Related Works’ section). This augmentation of dense intrinsic reward signals allows the agents to learn from the interactions that do not result in any extrinsic/environmental rewards.

### v2

# Enhancing Heterogeneous Multi-Agent Cooperation in Decentralized MARL via GNN-driven Intrinsic Rewards

@@ Abstract.
Key words and phrases:
Multi-agent Reinforcement Learning, Graph Neural Network, Intrinsic Rewards, Decentralized Training, Inter-agent Collaboration
1. Introduction
@@
The paradigm of Multi-agent Reinforcement Learning (MARL) is rapidly emerging to be pivotal in a broad spectrum of practical applications such as resource management (Ying and Dayong, 2005), autonomous vehicles (Cao et al., 2013), traffic signal control (Calvo and Dusparic, 2018a), supply chain management (Fuji et al., 2018), robotics (Lillicrap et al., 2019), robot swarms (Hüttenrauch et al., 2017), etc. These applications generally benefit from the efficient use of the diverse capabilities of heterogeneous agents. Moreover, the successful execution of tasks in these multi-agent systems requires the agents to adapt their behaviors to other agents for effective coordination rather than operating independently. The real-world deployment of these MARL systems typically involves the agents relying solely on the local environmental information and learning policies with infrequent environmental rewards (Wiewiora, 2010; Wakilpoor et al., 2020).
@@
Applications such as package transport (Gerkey and Mataric, 2002), traffic lights control (Calvo and Dusparic, 2018b), disaster response (Michael et al., 2012), agriculture (Ju and Son, 2019), etc. utilize agent heterogeneity such as distinct physical and behavioral traits of agents. Heterogeneity is also vital in multi-robot tasks as it enables efficient characterization and discovery of diverse behaviors, improving learning performance (Liu et al., 2022). On the other hand, the dependency on reward signals for the agent’s learning process introduces the issue of reward sparsity (Hare, 2019). Due to the lack of frequent feedback from the environment and the non-trivial nature of manually designing reward functions, MARL systems need to be robust enough to deal with infrequent environmental rewards.
Figure 1. Overview of the CoHet intrinsic reward architecture: utilizing the observation predictions of neighboring agents, it augments the self-supervised intrinsic rewards with the sparse environmental rewards to elicit collaborative actions
[…]
The figure gives an overview of the CoHet intrinsic reward calculation architecture – it shows how next observation predictions are utilized to calculate intrinsic rewards that are augmented with the extrinsic rewards.
@@
In addition to these challenges, the majority of real-world applications constrain the agents to act in a decentralized manner, and under partial observability, where each agent has a partial view of the shared environment. As a result, it is impractical for them to learn cooperative behaviors by utilizing a centralized algorithm that possesses global knowledge of all the agents and the state space (Iqbal and Sha, 2019; Liu et al., 2019). In comparison to centralized training and full observability, the challenges of agent heterogeneity and reward sparsity are more severe in decentralized training under partial observability (Ma et al., 2022). Despite real-world requirements, existing solutions often rely on global parameter sharing or a centralized critic.
@@
To the best of our knowledge, no prior research has addressed the issue of cooperative heterogeneous MARL in decentralized training settings under the practical constraints of real-world applications, such as partial observability and reward sparsity (see ‘Related Works’ section for more details). We propose CoHet, an algorithm that facilitates heterogeneous agent cooperation addressing the constraints required for real-world deployments. CoHet does not require any prior knowledge of agent heterogeneity (e.g speed, size, type, agent index). It employs an architecture for learning heterogeneous MARL policies by utilizing a novel Graph Neural Network-based intrinsic motivation/reward calculation mechanism. In summary, our specific contributions are as follows:
@@
A Novel Intrinsic Reward Mechanism: We introduce a novel self-supervised intrinsic reward calculation algorithm — CoHet [^1] The codebase for CoHet can be found at: https://github.com/jahirsadik/CoHet-Implementation, utilizing the underlying communication graph of a Graph Neural Network (GNN). In comparison to previous methods, CoHet accurately estimates the intrinsic rewards in the presence of agent heterogeneity (e.g. physical attributes/composition, behavioral differences) by using only the agents’ local neighborhood information. We present two formulations of the CoHet algorithm, one that utilizes the neighborhood predictions (CoHetteam), and another that uses the agent’s own predictions (CoHetself) for intrinsic reward calculation.
@@
Integration with Established Algorithms: Our standalone intrinsic motivation architecture can be integrated with existing decentralized heterogeneous policy learning algorithms, thus enhancing performance in cooperative MARL benchmarks. We demonstrate this by incorporating the state-of-the-art HetGPPO algorithm (Bettini et al., 2023a), leveraging its underlying GNN communication graph for intrinsic reward calculation. In contrast to previous heterogeneous policy learning techniques, this formulation requires no prior knowledge of the types of agent heterogeneity, indexing, etc.
@@
Extensive Validation and Scalability: We validate CoHet in the presence of heterogeneous agents in six different scenarios in the Multi-agent Particle Environment (MPE) and Vectorized Multi-Agent Simulator (VMAS) benchmarks, showing superior performance. We present findings on the impact of agent dynamics models on the intrinsic reward calculation, compare the two variants of the algorithm, and demonstrate its robustness to an increasing number of heterogeneous agents in a shared environment.
The subsequent sections will establish the necessary foundation of our work and explore our contributions in detail. In Section 2, we present the related works in this domain and highlight the lack of studies addressing real-world constraints. In Section 3, we formulate our problem and discuss the background necessary for our proposed method. Our proposed algorithm and system architecture are presented in Section 4. The experimental outcome of our research along with further studies has been delineated in Section 5. Finally, we conclude with an overview of our research and discuss the scope of potential future work in Section 6.
2. Related Works
@@
In response to the challenges posed by agent heterogeneity and the sparsity of rewards, most existing literature addresses one or the other. Existing methods addressing agent heterogeneity either necessitate prior knowledge of the types of heterogeneous agents (Foerster et al., 2016; Gupta et al., 2017) or suggest solutions that are specific to problems within only certain sub-classes of heterogeneity (Wang et al., 2020; Li et al., 2021; Lowe et al., 2020). Moreover, their use in partially observable systems is limited due to the lack of inter-agent communication. A notable study on heterogeneous MARL Bettini et al. (2023a), proposes an algorithm termed HetGPPO (Heterogeneous GNN-based Proximal Policy Optimization), that is capable of learning heterogeneous decentralized policies in partially observable scenarios. It uses a GNN-based communication layer for sharing information among agents within local neighborhoods, thereby mitigating the effects of partial observability. Unlike previous methods, their reliance on only local information enables Decentralized Training with Decentralized Execution (DTDE). Similar to their work, CoHet utilizes only the local neighborhood information under the constraints of DTDE.
@@
Although the aforementioned studies address the problem of agent heterogeneity, typically involving centralized critics, parameter sharing, or prior knowledge of agent heterogeneity, they do not consider inter-agent cooperation under reward sparsity. Dense scalar reward signals termed “Intrinsic Motivation”, are often used to encourage exploration or coordination among agents (Du et al., 2019; Jaques et al., 2019; Jeon et al., 2022). In case of the algorithm proposed by Ma et al. (2022) termed ELIGN (Expectation Alignment as a Multi-Agent Intrinsic Reward), the intrinsic rewards foster inter-agent coordination in a decentralized manner. However, a significant drawback of using an agent’s own dynamics model as a proxy to calculate neighborhood predictions in ELIGN is that it becomes more challenging for agents to accurately model the dynamics of other agents in the presence of agent heterogeneity. Inaccuracies in the dynamics model can result in misleading alignment signals, and as a result, ELIGN scales poorly in the presence of heterogeneous agents. In contrast, CoHet utilizes the local neighborhood information passed via the underlying GNN-communication graph, to more accurately model the heterogeneity among agents.
@@
There are only a couple of existing research works that simultaneously tackle agent heterogeneity under constraints of partial observability and reward sparsity (Zheng et al., 2020; Andres et al., 2022). In the former work, heterogeneity is defined differently, referring to a mixture of on-policy, off-policy, and Evolutionary Algorithm (EA) agents and not the diverse physical and behavioral traits of agents. Moreover, their adoption of a local-global memory replay prevents them from undergoing training in a fully decentralized manner. The latter leverages intrinsic motivation to tackle heterogeneous agent cooperation under reward sparsity similar to our work. However, their utilization of a centralized critic that merges all agent parameters into a single network constrains its applicability in the DTDE paradigm. Despite its significance, the lack of solutions in decentralized training settings limits the deployment of MARL agents in practical applications. CoHet addresses this notable research gap in the area of cooperative heterogeneous MARL, by fostering cooperation in a decentralized manner, under the real-world challenges of partial observability and reward sparsity.
3. Background
@@
This section outlines the Markov games framework used to formulate our problem and the message-passing Graph Neural Network employed for differentiable inter-agent communication of local observations and predictions.
3.1. Markov Games
@@
The Markov games framework (Littman, 1994) is a generalization of a Markov decision process to the case of multiple agents with cooperating or competing goals. The Partially Observable Markov Games (POMG) framework is used under conditions of partial observability. It is defined using the tuple —
@@
where, $\mathcal{V}=\{1,2,\ldots,N\}$ represents the set of all $N$ agents. The state space is denoted by $\mathcal{S}$. The observation space is defined as $\mathcal{O}\equiv\mathcal{O}_{1}\times\mathcal{O}_{2}\times\ldots\times\mathcal{O}_{n}$, where $\mathcal{O}_{i}\subseteq\mathcal{S}$ for each agent $i$ in $\mathcal{V}$. The action space for all agents is denoted by $\mathcal{A}\equiv\mathcal{A}_{1}\times\mathcal{A}_{2}\times\ldots\times\mathcal{A}_{n}$. Each agent $i\in\mathcal{V}$ has an observation instance $o_{i}\in\mathcal{O}_{i}$, which represents a partial view of $\mathcal{S}$. The reward function for agent $i$ is given by $\mathcal{R}_{i}\colon\mathcal{S}\times\mathcal{A}\times\mathcal{S}\mapsto\mathcal{R}$. The stochastic state transition model is denoted by $\mathcal{T}\colon\mathcal{S}\times\mathcal{A}\times\mathcal{S}\mapsto[0,1]$, describing how each agent transitions to the next state, given the current state and the action taken. Finally, $\gamma$ represents the discount factor, where $0\leq\gamma\leq 1$. Additionally, the reward that agent $i$ receives at time step $t$ can be denoted using $r_{i}^{t}=\mathcal{R}_{i}(s^{t},a^{t},s^{t+1})$, where $a^{t}=(a_{1}^{t},a_{2}^{t},\ldots,a_{n}^{t})$ is the joint action taken by all the agents at time step $t$. The goal of each agent is to maximize the total expected discounted return $R_{i}=\sum^{T}_{t=0}\gamma^{t}r_{i}^{t}$ over the course of an episode with horizon $T$.
3.2. Message-passing Graph Neural Network
The message-passing technique (Kipf and Welling, 2017) used by GNNs to transfer information from one node to another has proven to be an effective learning framework for understanding the patterns, the neighborhood of nodes, and the sub-graphs in large graphs (Zhou et al., 2021). In our problem formulation, we define graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where
@@ 3.2. Message-passing Graph Neural Network
For the message-passing step in our GNN, at each time step, the agent embeddings and edge features are first computed and then utilized in the message-passing GNN kernel to learn the local sub-graph. Information is iteratively passed between adjacent agents along the edges of the graph structure. The use of only the non-absolute observation features as agent embeddings allows the outputs of the message-passing GNN kernel to be invariant to geometric translations, thereby enhancing generalization. In order to incorporate the CoHet architecture on top of the underlying GNN formulation of HetGPPO, the inputs to our message-passing GNN kernel consist of the agent embedding $z_{i}=\omega_{\theta_{i}}(x_{i})$ and edge attributes $e_{ij}=p_{ij}\|v_{ij}$. Here, $\omega_{\theta_{i}}$ represents a Multi-Layer Perceptron (MLP) encoder with parameters $\theta_{i}$, and $x_{i}$ represents the non-absolute features of the agent obtained by removing absolute geometric features from the observation $o_{i}$. Edge attributes $e_{ij}$ are calculated using the absolute features of agent observations, such as position $p_{ij}$ and relative velocity $v_{ij}$. Using local information from all neighbors $j\in\mathcal{N}_{i}$, the GNN model output $h_{i}$ for agent $i$ is calculated in Equation 1.
@@
In Equation 1, $\psi_{\theta_{i}}$ and $\phi_{\theta_{i}}$ are two MLPs parameterized by $\theta_{i}$ and the aggregation operator $\bigoplus$ sums the $\phi_{\theta_{i}}$ outputs for all the neighbors of agent $i$. Finally, two distinct MLP decoders take the GNN output $h_{i}$ and produce the value $V_{i}(o_{\mathcal{N}_{i}})$ and the action $a_{i}$, distributed according to $a_{i}\sim\pi_{i}(\cdot\mid o_{\mathcal{N}_{i}})$. This formulation of GNN allows us to utilize it for both intrinsic reward calculation and heterogeneous policy learning of HetGPPO, based on local neighborhood information.
@@
29:         $r^{t}\leftarrow r^{t}_{total_{1}}\|r^{t}_{total_{2}}\|\ldots\|r^{t}_{total_{N}}$
30:         $\mathcal{D}_{k}\leftarrow\mathcal{D}_{k}\cup(o^{t},a^{t},r^{t},o^{t+1})$
@@
4. The CoHet Algorithm
Real-world deployment of multi-agent systems requires agents that can deal with challenges such as decentralized training, operating with a partial view of the environment, and learning from infrequent environmental feedback signals. In this section, we introduce CoHet, a decentralized algorithm designed to enhance cooperation among heterogeneous agents in partially observable environments with sparse rewards. It provides a standalone self-supervised intrinsic reward architecture that can be incorporated with existing decentralized policy optimization algorithms. It fosters the learning of collaborative behaviors by reducing future uncertainty within each agent’s neighborhood. CoHet encourages the agents to align their actions with their neighbors’ predictions by imposing intrinsic reward penalties that deter deviations from such alignment. Furthermore, these calculated rewards serve as a source of dense reward signals that facilitate policy learning in numerous real-world tasks where manually designing reward functions is infeasible.
4.1. Algorithm Description